# Intent-Based Access Control: Using LLMs to Intelligently Manage Access Control

### v2

@@ Abstract.
In every enterprise database, administrators must define an access control policy that specifies which users have access to which assets. Access control straddles two worlds: policy (organization-level principles that define who “should” have access) and process (database-level primitives that actually implement the policy). Assessing and enforcing process compliance with a policy is a manual and ad-hoc task. This paper introduces a new paradigm for access control called Intent-Based Access Control for Databases (IBAC-DB). In IBAC-DB, access control policies are expressed more precisely using a novel format, the natural language access control matrix (NLACM). Database access control primitives are synthesized automatically from these NLACMs. These primitives can be used to generate new DB configurations and/or evaluate existing ones. This paper presents a reference architecture for an IBAC-DB interface, an initial implementation for PostgreSQL (which we call LLM4AC), and initial benchmarks that evaluate the accuracy and scope of such a system. We find that our chosen implementation, LLM4AC, vastly outperforms other baselines, achieving near-perfect F1 scores on our initial benchmarks.
@@ 7.4. Differencer Results
8. DDepSys: An Initial System
@@ 8.2. Dependency Resolver
The dependency resolver takes as input backend $B$, policy tables $P$, policy documentation $P_{Doc}$, and requirement labels generated during requirement identification $L_{R}$. The output is an execution graph $G=(V,E)$ where nodes are policy tables and edges represent required inputs. We use Chain-of-Thought prompting, following the intuition that we are really attempting to generate instructions for implementing deployment requirements on $B$ using the NL in $P$ and $P_{Doc}$: You are a database security analyst. Your job is to make sure excessive access privileges are not granted to anyone. Consider the following tables representing deployment policies: $L_{R}$, $P$. Consider the following documentation for these tables: $P_{Doc}$. Write a set of instructions for implementing these policies on a $B$ backend. Once we have instructions, then we ask followup questions to generate $G$: For each step in the instructions, which policy tables from $L_{R}$ are needed? and For each step in the instructions, list the previous steps whose outputs will be needed for this step.
@@ 8.3. Primitive Implementation
Using the given requirement labels and execution graph, DDepSys uses a set of general and extensible primitives to synthesize and audit deployment policy implementation. DDepSys achieves this using the following templated prompts for synthesis and auditing.
@@ 8.3.2. Auditing Primitives
Differencing Once such implementation tables $I$ have been generated, they must be differenced to ensure they comply with policy $P$. Given policy table and corresponding implementation table $P_{i}$ and $I_{i}$, we determine which privileges are in implementation that are not permitted according to policy in two steps: (i) deployment element identification DDepSys must first extract the types of database elements needed to define a given type of deployment requirement (we call this set of types a deployment element). For example, in a table describing role hierarchies with a ”Role” and ”Child” column, the deployment element is an inheritance relationship, consisting of both parent and child. (ii) deployment element matching: determine which database elements (e.g., roles/users, tables/views) appear in both $P_{i}$ and $I_{i}$ (iii) privilege comparison: for database element matches, determine whether implementation privileges in $I_{i}$ are permitted according to privileges in $P_{i}$. If not, then $I_{i}$ does not comply with $P_{i}$, and DDepSys records the mismatching privilege.
@@ 8.4. Synthesis Feedback
9. DDepSys Evaluation
In this section, we discuss how to evaluate a system for synthesizing and auditing formal specification such as DDepSys. We develop an initial macrobenchmark and detailed microbenchmark for evaluating DDepSys with respect to synthesis and auditing accuracy. Then, we provide a detailed design for a more in-depth evaluation of DDepSys.
Evaluation Challenges. As mentioned in Section 1, DDepSys will be used to reduce the human effort required in developing formal specifications from policy documents themselves. As a first step toward evaluating whether DDepSys reduces human effort, we view system accuracy as a proxy for human effort, as the fewer errors a human has to detect and correct, the less effort they will require. Further, since DDepSys is the first generalized automated system to synthesize and audit deployment policy, there is no existing system to evaluate DDepSys against in a macrobenchmark. As a first step toward a macrobenchmark, we compare the system accuracy of DDepSys against a task-specific variant. We then discuss a full design for a macrobenchmark involving a user study.
We evaluate DDepSys on a realistic benchmark and find that DDepSys performs almost as well as a task-specific alternatives, and that DDepSys’s performance is not affected by choice of input and metadata formats. However, we also observe that DDepSys’s performance is limited due to current limitations of LLMs, giving rise to novel research challenges in using LLMs for deployment.
@@ 9.1. Benchmark Details
We evaluate DDepSys with respect to access control deployment requirements: access control privileges, role hierarchies, and temporal access control. This means the benchmark requires policy tables consisting of: (i) a role hierarchy (ii) Descriptions of database views (iii) descriptions of allowed SQL operators (iv) descriptions of temporal conditions, e.g., ”9 to 5 on weekdays” (v) the ground truth implementations for each of the above on all backends we will test. We describe them in detail here.
@@
Generating Roles & Views We use database views derived from NL questions representing SQL queries from the recent BIRD benchmark. These queries are over the european_football_2 database, which is the database with the largest schema in the benchmark, with 8 tables and 201 columns. Note that we use the database with the largest schema (as opposed to, say, database size) because schema size affects the size of context needed as input to DDepSys, which affects performance. We use roles from a real-world role hierarchy of the Amazon Access dataset (cite amazon access). As the roles in this dataset are anonymized, we use GPT-4 to generate NL descriptions of roles for a corporate role hierarchy. These generated role and view descriptions are used in the cells of the ’Role’ column and as the column headers of the NLACM and TACM, and we use the hierarchy tree given by the Amazon Access dataset to organize the generated role descriptions into ’Role’ and ’Child’ columns in the RHL.
@@
Implementation Tables for Differencing To test auditing (DDepSys’s differencing procedure), the benchmark also must include implementation tables containing the ground truth database deployment code for each policy table. We use the ground truth SQL from the BIRD benchmark for database views, the role names matching each description for database roles, and we manually generate implementation for the access control privileges, role hierarchy, and temporal access control conditions.
@@
9.2. Macrobenchmark: DDepSys vs Task-Specific
We compare DDepSys against a task-specific version of DDepSys, which uses prompts and synthesis and auditing procedures specific for implementing access control privileges, temporal access control, and role hierarchies.
Synthesis. DDepSys correctly generates 194 / 285. The task-specific variant correctly generates 243 / 285. However, DDepSys has fewer failure modes, which may make it easier for the end user to debug. We analyze the errors more closely: DDepSys has view errors due to deployment element mismatches reintroduced by the feedback loop (38), incorrect privilege-inheritance generation (12), and incorrect temporal constraint generation due to stray GRANT statements (41). On the other hand, the task-specific variant has the same failure modes, and more: view errors due to PK/FK misunderstanding or column values (19), incorrect privilege-inheritance generation (3), incorrect temporal constraint generation due to stray GRANT statements (5), or due to incorrect temporal constraint specification (15).
Auditing. DDepSys’s novel differencing procedure has a F1 score of 0.71. The task-specific variant has a F1 score of 0.85. DDepSys’s sources of error are false positive role inheritance matches, false negative temporal constraint matches and minor feedback-related LLM hallucinations The task-specific variant’s sources of error are false positive role inheritance matches.
@@ 9.3. Microbenchmarks
The reference architecture for DDepSys has a large design space characterized by LLM choice, database backends, and schema and metadata formats. We evaluate DDepSys with respect to each of these design space aspects, and we find that many LLMs simply lack the knowledge to generate deployment requirement implementations.
@@ 9.3.1. LLM Choice
We evaluate DDepSysusing closed, small open, and large specialized open LLMs. We use gpt-3.5-turbo (ChatGPT) as the closed LLM, mistral-7b-instructv0.1 as the small open LLM, and CodeLlama-70B as the large specialized open LLM.
@@ 9.3.2. Database Backend
We evaluate DDepSys with respect to two backends with different SQL variants for access control implementation (Postgres and MySQL) and one non-SQL backend (MongoDB).
Synthesis. DDepSys on Postgres has an accuracy of 194 / 285. On MySQL, the accuracy was 180 / 285–the results are exactly the same, but for one temporal condition, ChatGPT does not realize it cannot use a CREATE POLICY statement, as MySQL does not support row-level security directly. On the other hand, MongoDB has an accuracy of 42 / 285. The high number of errors is due to major python parsing issues and errors, and ChatGPT is completely unable to implement temporal constraints.
@@
In summary, DDepSys’s performance depends on the LLM’s knowledge of database backend semantics, but LLMs’ knowledge of certain backends, like MongoDB, is currently limited.
@@ 9.3.3. Schema Format
We evaluate DDepSys with respect to three different formats for representing the database schema and data dictionary metadata from the catalog to ChatGPT: as JSON objects where keys are table names and values are columns, as tables with a ’Table’ column and a ’Column’ column, and as text documents.
Synthesis. DDepSys on JSON achieves accuracy 194 / 285. DDepSys on Tabular achieves accuracy 189 / 285. Here, the results are exactly the same, except for 5 hallucinations arbitrary hallucinations. DDepSys on Unstructured Text achieves accuracy 191 / 285. Here also, the results are exactly the same, except for 3 arbitrary hallucinations–distinct from the hallucinations of Tabular.
Auditing. DDepSys on JSON achieves an F1 score of 0.71. DDepSys on Tabular achieves an F1 score of 0.65. Here, the results are exactly the same, except for 9 arbitrary hallucinations. DDepSys on Unstructured Text achieves 0.68. Here, the results are exactly the same, except for 4 arbitrary hallucinations.
In summary, we observe that varying the format in which database schemas and data dictionaries from catalogs are shown to the LLM has little effect on DDepSys’s performance.
@@ 9.3.4. Policy Format
We evaluate DDepSys with respect to two different formats for representing the policy: the tabular representations we have proposed, such as NLACMs, and documents consisting of sentences where each sentence represents a table cell, describing either permitted SQL operations, temporal constraints, or a role hierarchy relationship.
Synthesis. DDepSys on tables achieves accuracy 194 / 285. DDepSys on documents achieves accuracy 187 / 285. Here, the results are exactly the same, except for 7 extra GRANT statements.
Auditing. DDepSys on tables achieves an F1 score of 0.71. DDepSys on documents achieves an F1 score of 0.65. Here, the results are exactly the same, except for 9 arbitrary hallucinations.
In summary, we observe that varying the format in which deployment policies shown to the LLM has little effect on DDepSys’s performance. While documents may underperform slightly compared to tables, we conjecture this only occurs because representing documents as a string to LLMs requires more tokens than tables, and LLMs currently cannot perform accurately with large volumes of context. The difference in performance with respect to policy formats will likely diminish as LLMs improve.
@@ 10. Conclusion
In this paper, we recognize the problem of automating the policy comparison and auditing of access control policies written in NL. To facilitate this, we propose IBAC-DB, a new access control paradigm. We define a reference architecture for IBAC-DB which can be applied across several different LLMs and DBMSs. Then, we build one possible system, LLM4AC, using this reference architecture. We find that LLM4AC performs well compared to other role-view mapping methods, and that LLM4AC performs reliably: it has low performance variation by database and NL perturbation. We build systems to handle other database deployment requirements as well, RHieSys and DDepSys, and we find promising initial results.

### v3

@@ Abstract.
In every enterprise database, administrators must define an access control policy that specifies which users have access to which assets. Access control straddles two worlds: policy (organization-level principles that define who “should” have access) and process (database-level primitives that actually implement the policy). Assessing and enforcing process compliance with a policy is a manual and ad-hoc task. This paper introduces a new paradigm for access control called Intent-Based Access Control for Databases (IBAC-DB). In IBAC-DB, access control policies are expressed more precisely using a novel format, the natural language access control matrix (NLACM). Database access control primitives are synthesized automatically from these NLACMs. These primitives can be used to generate new DB configurations and/or evaluate existing ones. This paper presents a reference architecture for an IBAC-DB interface, an initial implementation for PostgreSQL (which we call LLM4AC), and initial benchmarks that evaluate the accuracy and scope of such a system. We further describe how to extend LLM4AC to handle other types of database deployment requirements, including temporal constraints and role hierarchies. We propose RHieSys, a requirement-specific method of extending LLM4AC, and DePLOI, a generalized method of extending LLM4AC. We find that our chosen implementation, LLM4AC, vastly outperforms other baselines, achieving high accuracies and F1 scores on our initial Dr. Spider benchmark. On all systems, we find overall high performance on expanded benchmarks, which include state-of-the-art NL2SQL data requiring external knowledge, and real-world role hierarchies from the Amazon Access dataset.
@@ 7.4. Differencer Results
8. DePLOI: An Initial System
@@ 8.2. Dependency Resolver
The dependency resolver takes as input backend $B$, policy tables $P$, policy documentation $P_{Doc}$, and requirement labels generated during requirement identification $L_{R}$. The output is an execution graph $G=(V,E)$ where nodes are policy tables and edges represent required inputs. We implement and test the dependency resolver in two ways: (i) using a prompt that generates the instructions $Inst$, (ii) using human instructions. We use Chain-of-Thought prompting as the prompt for dependency resolver, following the intuition that we are really attempting to generate instructions for implementing deployment requirements on $B$ using the NL in $P$ and $P_{Doc}$: You are a database security analyst. Your job is to make sure excessive access privileges are not granted to anyone. Consider the following tables representing deployment policies: $L_{R}$, $P$. Consider the following documentation for these tables: $P_{Doc}$. Write a set of instructions for implementing these policies on a $B$ backend. Once we have instructions, then we ask followup questions to generate $G$: For each step in the instructions, which policy tables from $L_{R}$ are needed? and For each step in the instructions, list the previous steps whose outputs will be needed for this step.
@@ 8.3. Primitive Implementation
Using the given requirement labels and execution graph, DePLOI uses a set of general and extensible primitives to synthesize and audit deployment policy implementation. DePLOI achieves this using the following templated prompts for synthesis and auditing.
@@ 8.3.2. Auditing Primitives
Differencing Once such implementation tables $I$ have been generated, they must be differenced to ensure they comply with policy $P$. Given policy table and corresponding implementation table $P_{i}$ and $I_{i}$, we determine which privileges are in implementation that are not permitted according to policy in two steps: (i) deployment element identification DePLOI must first extract the types of database elements needed to define a given type of deployment requirement (we call this set of types a deployment element). For example, in a table describing role hierarchies with a ”Role” and ”Child” column, the deployment element is an inheritance relationship, consisting of both parent and child. (ii) deployment element matching: determine which database elements (e.g., roles/users, tables/views) appear in both $P_{i}$ and $I_{i}$ (iii) privilege comparison: for database element matches, determine whether implementation privileges in $I_{i}$ are permitted according to privileges in $P_{i}$. If not, then $I_{i}$ does not comply with $P_{i}$, and DePLOI records the mismatching privilege.
@@ 8.4. Synthesis Feedback
9. DePLOI Evaluation
In this section, we discuss how to evaluate a system for synthesizing and auditing formal specification such as DePLOI. We develop an initial macrobenchmark and detailed microbenchmark for evaluating DePLOI with respect to synthesis and auditing accuracy. Then, we provide a detailed design for a more in-depth evaluation of DePLOI.
Evaluation Challenges. As mentioned in Section 1, DePLOI will be used to reduce the human effort required in developing formal specifications from policy documents themselves. As a first step toward evaluating whether DePLOI reduces human effort, we view system accuracy as a proxy for human effort, as the fewer errors a human has to detect and correct, the less effort they will require. Further, since DePLOI is the first generalized automated system to synthesize and audit deployment policy, there is no existing system to evaluate DePLOI against in a macrobenchmark. As a first step toward a macrobenchmark, we compare the system accuracy of DePLOI against a task-specific variant. We then discuss a full design for a macrobenchmark involving a user study.
We evaluate DePLOI on a realistic benchmark and find that DePLOI performs almost as well as a task-specific alternatives, and that DePLOI’s performance is not affected by choice of input and metadata formats. However, we also observe that DePLOI’s performance is limited due to current limitations of LLMs, giving rise to novel research challenges in using LLMs for deployment.
@@ 9.1. Benchmark Details
We evaluate DePLOI with respect to access control deployment requirements: access control privileges, role hierarchies, and temporal access control. This means the benchmark requires policy tables consisting of: (i) a role hierarchy (ii) Descriptions of database views (iii) descriptions of allowed SQL operators (iv) descriptions of temporal conditions, e.g., ”9 to 5 on weekdays” (v) the ground truth implementations for each of the above on all backends we will test. We describe them in detail here.
@@
Generating Roles & Views We use database views derived from NL questions representing SQL queries from the recent BIRD benchmark. These queries are over the european_football_2 database, which is the database with the largest schema in the benchmark, with 8 tables and 201 columns. Note that we use the database with the largest schema (as opposed to, say, database size) because schema size affects the size of context needed as input to DePLOI, which affects performance. We use roles from a real-world role hierarchy of the Amazon Access dataset (cite amazon access). As the roles in this dataset are anonymized, we use GPT-4 to generate NL descriptions of roles for a corporate role hierarchy. These generated role and view descriptions are used in the cells of the ’Role’ column and as the column headers of the NLACM and TACM, and we use the hierarchy tree given by the Amazon Access dataset to organize the generated role descriptions into ’Role’ and ’Child’ columns in the RHL.
@@
Implementation Tables for Differencing To test auditing (DePLOI’s differencing procedure), the benchmark also must include implementation tables containing the ground truth database deployment code for each policy table. We use the ground truth SQL from the BIRD benchmark for database views, the role names matching each description for database roles, and we manually generate implementation for the access control privileges, role hierarchy, and temporal access control conditions.
@@
9.2. Macrobenchmark: DePLOI vs Task-Specific
We compare DePLOI against a task-specific version of DePLOI, which uses prompts and synthesis and auditing procedures specific for implementing access control privileges, temporal access control, and role hierarchies.
Synthesis. DePLOI correctly generates 194 / 285. The task-specific variant correctly generates 243 / 285. However, DePLOI has fewer failure modes, which may make it easier for the end user to debug. We analyze the errors more closely: DePLOI has view errors due to deployment element mismatches reintroduced by the feedback loop (38), incorrect privilege-inheritance generation (12), and incorrect temporal constraint generation due to stray GRANT statements (41). On the other hand, the task-specific variant has the same failure modes, and more: view errors due to PK/FK misunderstanding or column values (19), incorrect privilege-inheritance generation (3), incorrect temporal constraint generation due to stray GRANT statements (5), or due to incorrect temporal constraint specification (15).
Auditing. DePLOI’s novel differencing procedure has a F1 score of 0.71. The task-specific variant has a F1 score of 0.85. DePLOI’s sources of error are false positive role inheritance matches, false negative temporal constraint matches and minor feedback-related LLM hallucinations The task-specific variant’s sources of error are false positive role inheritance matches.
@@ 9.3. Microbenchmarks
The reference architecture for DePLOI has a large design space characterized by LLM choice, database backends, and schema and metadata formats. We evaluate DePLOI with respect to each of these design space aspects, and we find that many LLMs simply lack the knowledge to generate deployment requirement implementations.
@@ 9.3.1. LLM Choice
We evaluate DePLOIusing closed, small open, and large specialized open LLMs. We use gpt-3.5-turbo (ChatGPT) as the closed LLM, mistral-7b-instructv0.1 as the small open LLM, and CodeLlama-70B as the large specialized open LLM.
@@ 9.3.2. Database Backend
We evaluate DePLOI with respect to two backends with different SQL variants for access control implementation (Postgres and MySQL) and one non-SQL backend (MongoDB).
Synthesis. DePLOI on Postgres has an accuracy of 194 / 285. On MySQL, the accuracy was 180 / 285–the results are exactly the same, but for one temporal condition, ChatGPT does not realize it cannot use a CREATE POLICY statement, as MySQL does not support row-level security directly. On the other hand, MongoDB has an accuracy of 42 / 285. The high number of errors is due to major python parsing issues and errors, and ChatGPT is completely unable to implement temporal constraints.
@@
In summary, DePLOI’s performance depends on the LLM’s knowledge of database backend semantics, but LLMs’ knowledge of certain backends, like MongoDB, is currently limited.
@@ 9.3.3. Schema Format
We evaluate DePLOI with respect to three different formats for representing the database schema and data dictionary metadata from the catalog to ChatGPT: as JSON objects where keys are table names and values are columns, as tables with a ’Table’ column and a ’Column’ column, and as text documents.
Synthesis. DePLOI on JSON achieves accuracy 194 / 285. DePLOI on Tabular achieves accuracy 189 / 285. Here, the results are exactly the same, except for 5 hallucinations arbitrary hallucinations. DePLOI on Unstructured Text achieves accuracy 191 / 285. Here also, the results are exactly the same, except for 3 arbitrary hallucinations–distinct from the hallucinations of Tabular.
Auditing. DePLOI on JSON achieves an F1 score of 0.71. DePLOI on Tabular achieves an F1 score of 0.65. Here, the results are exactly the same, except for 9 arbitrary hallucinations. DePLOI on Unstructured Text achieves 0.68. Here, the results are exactly the same, except for 4 arbitrary hallucinations.
In summary, we observe that varying the format in which database schemas and data dictionaries from catalogs are shown to the LLM has little effect on DePLOI’s performance.
@@ 9.3.4. Policy Format
We evaluate DePLOI with respect to two different formats for representing the policy: the tabular representations we have proposed, such as NLACMs, and documents consisting of sentences where each sentence represents a table cell, describing either permitted SQL operations, temporal constraints, or a role hierarchy relationship.
Synthesis. DePLOI on tables achieves accuracy 194 / 285. DePLOI on documents achieves accuracy 187 / 285. Here, the results are exactly the same, except for 7 extra GRANT statements.
Auditing. DePLOI on tables achieves an F1 score of 0.71. DePLOI on documents achieves an F1 score of 0.65. Here, the results are exactly the same, except for 9 arbitrary hallucinations.
In summary, we observe that varying the format in which deployment policies shown to the LLM has little effect on DePLOI’s performance. While documents may underperform slightly compared to tables, we conjecture this only occurs because representing documents as a string to LLMs requires more tokens than tables, and LLMs currently cannot perform accurately with large volumes of context. The difference in performance with respect to policy formats will likely diminish as LLMs improve.
9.3.5. Can LLMs Handle Dependency Resolving?
DePLOI assumes instructions for executing code are given as documentation by the user. However, DePLOI would reduce synthesis and auditing effort even more if LLMs did not need such documentation to resolve code dependencies among different deployment requirements. We experiment with few-shot and Chain-of-Thought prompts on the more powerful GPT-4 to generate the output execution graphs, but we find that DePLOI severely underperforms because GPT-4 cannot reason about such dependencies. We report the highest performance we observed over various prompting strategies. The best strategy was Chain-of-Thought.
Synthesis. DePLOI with LLM-based dependency resolving achieves 0.432 compared to 0.681 using user-provided documentation. This is because in addition to correct CREATE POLICY statements to implement temporal constraints, DePLOI synthesizes incorrect GRANT statements which permit the same SQL operations without the required temporal constraints. Despite providing GPT-4 several few-shot examples of this reasoning, GPT-4 underperformed.
Auditing. DePLOI with LLM-based dependency resolving achieves 0.53 compared to 0.71 using user-provided documentation. Similarly to the failure mode for synthesis, this is because in addition to correct CREATE POLICY statements to implement temporal constraints, DePLOI falsely indicates errors due to missing GRANT statements which would incorrectly permit the same SQL operations without the required temporal constraints.
@@ 10. Conclusion
In this paper, we recognize the problem of automating the policy comparison and auditing of access control policies written in NL. To facilitate this, we propose IBAC-DB, a new access control paradigm. We define a reference architecture for IBAC-DB which can be applied across several different LLMs and DBMSs. Then, we build one possible system, LLM4AC, using this reference architecture. We find that LLM4AC performs well compared to other role-view mapping methods, and that LLM4AC performs reliably: it has low performance variation by database and NL perturbation. We build systems to handle other database deployment requirements as well, RHieSys and DePLOI, and we find promising initial results.
Appendix A DePLOI: Human Inputs
DePLOI differs from RHieSys and LLM4AC in that it uses human inputs, namely the dependency resolver instructions and the policy documentation. We provide the exact text of these instructions and documentation here.
A.1. Dependency Resolver Instructions
We show our dependency resolver instructions in Figure 11. Note that these instructions are agnostic to the input policy format.
Figure 11. Human Instructions given as input to DePLOI
A.2. Policy Documentation
The policy documentation describes the format in which the policy is shown, the deployment requirements represented, and their semantics. We provide the exact text of this documentation provided as input to DePLOI in Figure 12.
Figure 12. Policy Documentation given as input to DePLOI. Documentation for the table policy format on top, document policy format on bottom.